OpenStudy (anonymous):

PLEASE HELP A GIRL OUT The position of an object at time t is given by s(t) = -1 - 13t. Find the instantaneous velocity at t = 8 by finding the derivative.

OpenStudy (precal):

take the derivative using the power rule and then sub 8 for t

OpenStudy (precal):

s ' (8) is what you are looking for

OpenStudy (anonymous):

Okay but how do I write that out?

OpenStudy (precal):

do you know the power rule? It is the most basic rule for taking derivatives in calculus. It is the easiest one of all.

OpenStudy (precal):

y=3x^2 y ' = 6x

OpenStudy (anonymous):

oh okay. why did you use those numbers though?

OpenStudy (precal):

just to give you an example of power rule y = x^2 y ' = 2x

OpenStudy (anonymous):

Okay, but how do I find what x is?

OpenStudy (precal):

for your solution you take the derivative ie s ' (t) and then sub 8 for t

OpenStudy (precal):

example f(x) = x^2 and I want f ' (3) then f ' (x) = 2x f '(3) = 2(3) f'(3)=6

OpenStudy (precal):

do you see what your solution is?

OpenStudy (anonymous):

s′(t)=−13

OpenStudy (anonymous):

right?

OpenStudy (precal):

yes and guess what s ' (8)=-13 also, only because you don't have a variable to sub into

OpenStudy (anonymous):

wait so this s ' (8)=-13 is all? There aren't any other steps?

OpenStudy (anonymous):

@precal

OpenStudy (precal):

no because we don't have a variable to sub into

OpenStudy (anonymous):

so I can just write: s ' (t)=−13 s ' (8)=-13 or do I need to write it differently? @precal

OpenStudy (precal):

the first one represents the first derivative of your function the second one represents the derivative at that particular time in this case t=8

OpenStudy (precal):

go with the second one the first one is the velocity the second one is the velocity at that given time

OpenStudy (anonymous):

Thanks :)

OpenStudy (precal):

yw